A man and his brother have a combined age of 164 years. 46 years ago, the man was 7 times as old as his brother. How old is the man? How old is the brother?
Answer by scott8148(6628)

"A man and his brother have a combined age of 164 years" __ m+b=164 __ subtracting b __ m=164-b

"46 years ago, the man was 7 times as old as his brother" __ m-46=7(b-46) __ distributing __ m-46=7b-322

substituting __ (164-b)-46=7b-322 __ 118-b=7b-322 __ adding b+322 __ 440=8b __ dividing by 8 __ 55=b

substituting __ m=164-(55) __ m=109